# Computer Vision-based Adaptive Control for Back Exoskeleton Performance Optimization

### v1

@@ II-B3 Support Modulation
To modulate the exoskeleton support between low, medium, and strong, we use a simple parameter $k_{pyl}$ to scale the assistive torque provided to the hip joints within a lifting cycle. We find subject-specific values for $k_{pyl}^{min}$, $k_{pyl}^{max}$ during experiments as explained in section III-A to find the minimum and maximum peak value of assistance which the subject feels comfortable with while lifting a light weight ($5$ kg) and a heavy weight ($15$ kg) respectively. Then, we find the $k_{pyl}^{medium}$ as the value at $65\%$ of the $k_{pyl}^{min},k_{pyl}^{max}]$ range, according to the optimal trend found and motivated in section IV-A.
@@ II-C Hardware
The hardware used is a single-motor back-support exoskeleton (BSE) developed at the National University of Singapore (NUS) [5], shown in Fig. 2 (a). It features a differential series elastic actuator (D-SEA) powered by a $120W$ servo motor (RMD-L-5015) to deliver lifting assistance, with a nominal assistive torque of $30Nm$. The exoskeleton ensures precise force control and active backdrivability, allowing support during lifting while preserving walking autonomy. To enable the CVAC, we use an RGB camera (12MP) and an Arducam Time of Flight (ToF) sensor (240x180 @ 30fps, sub-2cm accuracy), with data acquisition handled by a Raspberry Pi 4B. Aligned with Industry 5.0 principles, we offload the system from the pipeline computational tasks by taking advantage of the cloud. A wireless connection links the exoskeleton to a remote computational station (Apple M3 chip, 8-core CPU, 10-core GPU, 24GB RAM) running the pipeline. After inference, outputs are sent back to the Raspberry Pi to modulate assistance in real-time, achieving an average $100ms$ inference time ($10fps$) while offloading deep learning tasks to the cloud.
@@ III Experiments
For both baseline and validation experiments, participants were eligible if they were adults aged 186518-65, with a height of 160200160-200 cm, capable of providing informed consent, and in good health to perform weightlifting tasks safely. The experiments took place at the BioRobotics Lab at the National University of Singapore (NUS). A total of 24 male healthy subjects (age: $24,29\pm 2,29$ yo; height: $176,04\pm 5,30$ cm; weight: $71,98\pm 8,23$ kg) voluntarily participated in the experiments, and none of the participants reported having a previous low back injury at the time of data collection. Twelve subjects participated in baseline tests, and 12 took part in validation tests, following recommendations for pilot studies to ensure feasibility, precision, and applicability to future research [37]. The experiment protocol was approved by the NUS Institutional Review Board (IRB) Committee with Reference Code LH-20-021, and experiments were conducted in accordance with the Research Compliance Policy on Human Subjects and Biomedical Research. Before the study began, the investigators explained the general purpose of the study and obtained verbal and written informed consent from the participants. During all experiments, participants were continuously monitored to ensure safety and avoid accidents. In all experiments, the BSE was utilized in active mode, with muscular activation recorded for the back Erector Spinae Iliocostalis (ESI) and Longissimus (ESL) muscles on both the right and left sides, as well as the leg Biceps Femoris (BF) and Rectus Femoris (RF) muscles. Muscle locations were identified following the SENIAM EMG Placement Guidelines [38]. To standardize data collection, Maximum Voluntary Contraction (MVC) exercises were performed for each muscle, enabling subsequent signal normalization during analysis. EMG signals were recorded using the Delsys Trigno Wireless System (Delsys Inc., Boston, MA, USA) at a $2150$ Hz sampling frequency. Data analysis adhered to best practices [39], including the application of a butterworth bandpass filter with low and high cut-off frequencies of $20$ Hz and $450$ Hz, respectively. Signals were rectified, and the root mean square (RMS) was calculated using a $200$ ms sliding time window. For baseline tests, where participants performed tasks in fixed positions, average and peak activation values were analyzed over the entire acquisition period. For validation tests involving walking and lifting, the analysis focused exclusively on lifting cycles. Weights of $5$, $10$, and $15$ kg, also referred to as light weight (LW), medium weight (MW) and heavy weight (HW), were selected for testing based on laboratory equipment availability and safety standards. Three control strategies (low, medium, and strong assistance levels) were used. Preliminary trials were conducted with each participant to calibrate low assistance (LA) and strong assistance (SA) levels based on subjective perception while lifting a $5$ kg and $15$ kg weight, respectively. For baseline tests, medium assistance (MA) was defined as 50% of the range between LA and SA. For validation tests, medium assistance was set at 65% of this range, as informed by baseline results discussed in Section IV-A.
@@ III-A Baseline Experiments
As explained in Section II-A, we conducted experiments with 12 subjects to define a two-dimensional optimization space, identifying the appropriate support level based on the payload, considering muscular activity reduction, perceived discomfort, and user preference as performance metrics. Since we classify weights and support levels into three discrete categories (low, medium, and high), there are nine potential combinations to test the exoskeleton’s performance. Additionally, three baseline conditions (LW, MW, and HW) without the exoskeleton bring the total to 12 conditions. To minimize testing time and subject fatigue, we used fractional factorial design [40], reducing the experiments to eight while maintaining statistical significance. This design samples the boundaries, corners, and midpoint of the optimization space and compares them to the no-exoskeleton baseline. For baseline experiments, the subject stands in a fixed position and repetitively lifts a box of varying weights per condition, first without the exoskeleton and then with different support levels. Each lifting cycle is synchronized to 35 beats per minute (bpm) using a metronome to standardize movement across subjects. A lifting activity consists of seven cycles, lasting a total of 30 seconds. A minimum rest time of 3 minutes is provided between activities to prevent fatigue or injury. Performance metrics were sampled as follows: during preliminary trials, users identified their preferred assistance level (light or strong) for low and heavy weights, yielding samples of $J_{prf}(A_{s}^{i},P_{yl}^{i})$. We measured muscular activation in back and leg muscles, as described in Section III, and compared it to non-exoskeleton conditions for the same weight. The percentage reduction in muscular activity provided samples of the EMG-ORF $J_{EMG}(A_{s}^{i},P_{yl}^{i})$. After each trial with the exoskeleton, subjects completed a questionnaire to evaluate perceived discomfort for each condition as outlined in Section II-A. The combined ratings from these questions, processed using Equation 3, provided samples for the perceived discomfort representation function $J_{dsc}(A_{s}^{i},P_{yl}^{i})$.
@@ III-B Pipeline Validation Experiments
For pipeline validation experiments, we designed a dynamic protocol to evaluate the payload estimation pipeline and the exoskeleton’s adaptation capabilities in a more variable environment. The subject starts from a fixed position and walks to three boxes placed in fixed positions ahead, containing $5$, $10$, or $15$ kg each, as shown in Figure 2 (b). To enhance generalization, we shuffle the positions and contents of the boxes before each test. The subject approaches each box, lifts it individually, and returns to the starting position. This process is repeated three times, ensuring each box is lifted three times. We performed this protocol under three conditions: without the exoskeleton, with static light support (preselected based on subject characteristics), and with adaptive pipeline control. This setup allows us to compare muscular activity across conditions for each lifting cycle. As validation metrics, first of all, we keep track during the experiment of the pipeline classification outputs and ground truth. In this context, because latency is an important concern, we only consider correct classifications happening before the lifting cycle starts. We also measure reductions in muscular activity targeting the same muscles of Section III-A, by comparing the no-exoskeleton condition with the static and adaptive control conditions. We collect user feedback using the same discomfort questionnaire described in Section II-A, and we gather user preferences between static and adaptive control. This approach enables us to sample the Validation Representation Functions (VRF) $J_{EMG}(A_{s}^{i},P_{yl}^{i})$, $J_{dsc}(A_{s}^{i},P_{yl}^{i})$, and $J_{prf}(A_{s}^{i},P_{yl}^{i})$ for each static/adaptive assistance (StA/AdA) and payload level pairing, creating a final validation space. Using this space, we assess the performance of adaptive control compared to static control.
@@ IV-A1 EMG-ORF
By analyzing the muscular activation data from the experiments in Section III-A, we derive the EMG-ORF $J_{EMG}(A_{s},P_{yl})$ shown in Figure 3 (I-a). These functions identify optimal performance areas based solely on muscular activity reduction. For instance, reducing back muscle activity consistently requires high assistance, achieving a $6.32\%$ mean and $16.92\%$ peak reduction for low weights, around $9.62\%$ mean and $15.00\%$ peak reduction for medium weights, and a $6.13\%$ mean and a $9.14\%$ peak reduction for high weights. However, high assistance for low weights significantly increases peak activation in leg muscles ($15.00\%$), which benefit more from medium or light support. The peak EMG-ORF for leg muscles indicates that transitioning diagonally from low-weight light-assistance to high-weight strong-assistance maintains relatively stable peak leg muscle activation while maximizing mean reduction ($5.00\%$ mean reduction for medium assistance with medium weight). Overall, considering all muscles together, medium to high assistance is preferable for mean reduction while medium to low for peak. However, given the adverse effect of high assistance on peak leg muscle activation, a better approach is to use low assistance for low weights and high assistance for high weights. This strategy balances maximizing back muscle activity reduction while minimizing peak activation in the legs.
@@ IV-A2 Discomfort & Preference-ORF
As outlined in Section II-A, the discomfort function was sampled by combining the ratings for each question, adjusting for the negative meaning of discomfort by flipping the ratings of positively worded questions. Using Equation 3, we obtained discomfort samples for different $(A_{s}^{i},P_{yl}^{i})$ pairs. For user preferences, we collected responses on LA/SA preference for LW/HW scenarios. By fitting these samples with GP, we derived continuous formulations of $J_{dsc}$ and $J_{prf}$ within the optimization space, as shown in Figure 3 (I-b). The results reveal that perceived discomfort peaks for medium weights with medium-to-high support. Overall, low assistance is preferable to minimize discomfort, regardless of the payload. Conversely, user preferences indicate a tendency to favor high assistance for heavy weights and low assistance for light weights.
@@ IV-A3 Total-ORF
It’s clear that different ORFs’ contributions can be considered individually to select the best assistance for a given payload for a specific metric optimization. In our case, to maximize acceptability and effectiveness, we combine the normalized versions of EMG-ORF, D-ORF, and P-ORF into a T-ORF $J_{tot}$, considering EMG contributions such as mean and peak reductions in back and overall muscle activity, as shown in Figure 3 (I-c). Results suggest that to optimize EMG reduction, discomfort, and user preference, assistance should be low for light weights and increase with the payload, following a non-linear trend. Based on this, we focus on optimizing “Mean All” activity to minimize muscle activation while avoiding excess leg muscle activation. To determine the optimal assistance for each payload, we compute the partial derivative of $J_{tot}$ with respect to assistance and slice the function along the payload axis. Setting the derivative to zero ($\frac{\partial J_{tot}}{\partial A_{s}}=f(P_{yl},A_{s})=0$) for each chosen $P_{yl}$ yields the optimal $A_{s}^{opt}$ corresponding to each selected $P_{yl}^{i}$, as shown in Figure 4. We then fit the points to an exponential function describing the relationship between payload and optimal support:
$P_{yl}=a*e^{b*A_{s}^{opt}}+c$ (6)
with $a=0.72$, $b=1.10$, and c=1.20c=-1.20. This function identifies an optimal region in the assistance-payload space, suggesting that if we had a way to continuously estimate payloads, we’d always be able to select the best-performing assistance in terms of our defined metrics. Following these findings, instead of the midpoint between low and high assistance levels we select 65% assistance during validation for medium weights. We provide in the supplementary materials the $a$, $b$, $c$ values found for different EMG contributions.

### v2

@@ II-B3 Support Modulation
To modulate the exoskeleton support between low, medium, and strong, we use a simple parameter $k_{pyl}$ to scale the low-level assistive torque profile provided to the hip joints within a lifting cycle, described in Section II-D. We find subject-specific values for $k_{pyl}^{min}$, $k_{pyl}^{max}$ during experiments as explained in section III-A to find the minimum and maximum peak value of assistance which the subject feels comfortable with while lifting a light weight ($5$ kg) and a heavy weight ($15$ kg) respectively. Then, we find the $k_{pyl}^{medium}$ as the value at $73\%$ of the $k_{pyl}^{min},k_{pyl}^{max}]$ range, according to the optimal trend found and motivated in section IV-A.
@@ II-C Hardware
The hardware we used is a single-motor back-support exoskeleton developed at the National University of Singapore (NUS) [11], shown in Fig. 2 (a). It features a differential series elastic actuator (D-SEA) powered by a $120W$ servo motor (RMD-L-5015) to deliver lifting assistance, with a nominal assistive torque of $30Nm$. The exoskeleton ensures precise force control and active backdrivability, allowing support during lifting while preserving walking autonomy. To enable the CVAC, we use an RGB camera (12 MP) and an Arducam Time-of-Flight (ToF) sensor (240×180 @ 30 fps, <<2 cm accuracy). We mount both sensors at the right hip (Figure 2 (b)), which provides a stable and minimally motion-affected vantage point. Data acquisition is handled by a Raspberry Pi 4B (RPI). Aligned with Industry 5.0 principles, we offload the system from the pipeline computational tasks by taking advantage of the cloud. A wireless connection links the exoskeleton to a remote computational station (Apple M3 chip, 8-core CPU, 10-core GPU, 24GB RAM) running the pipeline. After inference, results are returned to the RPI to modulate assistance in real time, as outlined in Figure 1. End-to-end, the CVAC pipeline achieves an average latency of 100 ms (10 fps) while offloading deep-learning computation to the cloud.
II-D Low-Level Control of Back-Support Exoskeleton
As shown in Figure 3, the mid-level assistance law computes the desired hip torque $\tau_{\mathrm{cmd}}(\theta,\dot{\theta},\ddot{\theta})$ as the sum of a quasi-static gravity-support term $k_{\theta}\theta$ and a direction-dependent velocity-shaping term gated by a smooth dynamic detector in the $(\dot{\theta},\ddot{\theta})$ plane:
$\tau_{\mathrm{cmd}}=k_{\theta}\,\theta-k_{v}\,\sigma\!\Big(\alpha_{e}\!\left[\left(\tfrac{\dot{\theta}}{\omega_{e}}\right)^{2}+\left(\tfrac{\ddot{\theta}}{a_{e}}\right)^{2}-1\right]\Big)\;\dot{\theta}\,\sqrt{\dot{\theta}^{2}+\varepsilon},$ (6)
where $\sigma(\cdot)$ denotes a sigmoidal activation function. Under the adopted sign convention ($\dot{\theta}>0$ during lowering and $\dot{\theta}<0$ during lifting), the velocity term attenuates gravity assistance during lowering and enhances it during lifting, with magnitude scaling approximately with $\dot{\theta}^{2}$. The resulting $\tau_{\mathrm{cmd}}$ is tracked by the low-level SEA torque controller (PID + disturbance observer), which closes the loop on measured hip torque. Payload adaptation is applied via $\tau_{\mathrm{adj}}=\tau_{\mathrm{cmd}}\cdot k_{\mathrm{pyl}}^{j}$.
[…]
Figure 3: Block diagram of the low-level control architecture. The mid-level assistance law combines gravity support and velocity shaping to generate $\tau_{\mathrm{cmd}}$. The CVAC-driven gain $k_{pyl}^{j}$ scales the desired torque, and the SEA torque controller (PID + DoB) compares the spring-deflection torque estimate with $\tau_{\mathrm{adj}}$ to command motor current and regulate the measured hip torque.
@@ III Experiments
For both baseline and validation experiments, participants were eligible if they were adults aged 186518-65, with a height of 160200160-200 cm, capable of providing informed consent, and in good health to perform weightlifting tasks safely. The experiments took place at the BioRobotics Lab at the National University of Singapore (NUS). A total of 24 male healthy subjects (age: $24,29\pm 2,29$ yo; height: $176,04\pm 5,30$ cm; weight: $71,98\pm 8,23$ kg) voluntarily participated in the experiments, and none of the participants reported having a previous low back injury at the time of data collection. Twelve subjects participated in baseline tests, and 12 took part in validation tests, following recommendations for pilot studies to ensure feasibility, precision, and applicability to future research [16]. The experiment protocol was approved by the NUS Institutional Review Board (IRB) Committee with Reference Code LH-20-021, and experiments were conducted in accordance with the Research Compliance Policy on Human Subjects and Biomedical Research. Before the study began, the investigators explained the general purpose of the study and obtained verbal and written informed consent from the participants. During all experiments, participants were continuously monitored to ensure safety and avoid accidents. In all experiments, the BSE was utilized in active mode, with muscular activation recorded for the back Erector Spinae Iliocostalis (ESI) and Longissimus (ESL) muscles on both the right and left sides, as well as the leg Biceps Femoris (BF) and Rectus Femoris (RF) muscles. We identified muscle locations following the SENIAM EMG Placement Guidelines [29]. To standardize data collection, Maximum Voluntary Contraction (MVC) exercises were performed for each muscle, enabling subsequent signal normalization during analysis. We recorded EMG signals using the Delsys Trigno Wireless System (Delsys Inc., Boston, MA, USA) at a $2150$ Hz sampling frequency. Data analysis adhered to best practices [18], including the application of a butterworth bandpass filter with low and high cut-off frequencies of $20$ Hz and $450$ Hz, respectively. We rectified signals and calculated the root mean square (RMS) using a $200$ ms sliding time window. For baseline tests, where participants performed tasks in fixed positions, we analyzed average and peak activation values over the entire acquisition period. For validation tests involving walking and lifting, we focused the analysis exclusively on lifting cycles. We selected weights of $5$, $10$, and $15$ kg, also referred to as light weight (LW), medium weight (MW) and heavy weight (HW), for testing based on laboratory equipment availability and safety standards. We used three control strategies (low, medium, and strong assistance levels). Given that the number of assistance-payload combinations increases quadratically with granularity, a finer discretization (e.g., five levels) would have substantially expanded the experimental matrix and prolonged the experimental sessions. For this reason, we limited the number of assistance levels to preserve data quality and avoid muscular and cognitive fatigue. The three-level assistance structure was intentionally selected as an experimental validation step to assess whether the continuous optimal assistance trend identified by our multi-metric analysis could be practically approximated and integrated with a pre-lift payload estimation pipeline. This coarser discretization also ensured reliability, as continuous estimation is sensitive to visual noise and small errors, especially at higher payloads. The 3-class design therefore prioritized stability and safety. The maximum equivalent torque at the hip ranged from 10 Nm to 30 Nm. We conducted preliminary trials with each participant to determine the low assistance (LA; $10,80\pm 0,98$ Nm) and strong assistance (SA; $24,00\pm 1,00$ Nm) torque levels. These values correspond to the mean $\pm$ standard deviation of the torques selected by the participants during the calibration based on subjective perception while lifting 5 kg and 15 kg loads, respectively, within the available torque range. For baseline tests, we defined medium assistance (MA) as 50% of the range between LA and SA. For validation tests, we set medium assistance at 73% of this range, as informed by baseline results discussed in Section IV-A.
@@ III-A Baseline Experiments
As explained in Section II-A, we conducted experiments with 12 subjects to define a two-dimensional optimization space, identifying the appropriate support level based on the payload, considering muscular activity reduction, perceived discomfort, and user preference as performance metrics. Since we classify weights and support levels into three discrete categories (low, medium, and high), there are nine potential combinations to test the exoskeleton’s performance. Additionally, three baseline conditions (LW, MW, and HW) without the exoskeleton bring the total to 12 conditions. To further minimize testing time and subject fatigue, we used fractional factorial design [3], reducing the experiments to eight while maintaining statistical significance. This design samples the boundaries, corners, and midpoint of the optimization space and compares them to the no-exoskeleton baseline. For baseline experiments, the subject stands in a fixed position and repetitively lifts a box of varying weights per condition, first without the exoskeleton and then with different support levels. Each lifting cycle is synchronized to 35 beats per minute (bpm) using a metronome. Participants were instructed to perform stoop lifting activities, lowering on every odd beep and lifting on every even beep to standardize movement speed and frequency across subjects. A lifting activity consists of seven cycles, lasting a total of 30 seconds. A minimum rest time of 3 minutes is provided between activities to prevent fatigue or injury. We sampled performance metrics as follows: during preliminary trials, users identified their preferred assistance level (light or strong) for low and heavy weights, yielding samples of $J_{prf}(A_{s}^{i},P_{yl}^{i})$. We measured muscular activation in back and leg muscles, as described in Section III, and compared it to non-exoskeleton conditions for the same weight. The percentage reduction in muscular activity provided samples of the EMG-ORF $J_{EMG}(A_{s}^{i},P_{yl}^{i})$. After each trial with the exoskeleton, subjects completed a questionnaire to evaluate perceived discomfort for each condition as outlined in Section II-A. The combined ratings from these questions, processed using Equation 3, provided samples for the perceived discomfort representation function $J_{dsc}(A_{s}^{i},P_{yl}^{i})$.
@@ III-B Pipeline Validation Experiments
For pipeline validation experiments, we designed a dynamic protocol to evaluate the payload estimation pipeline and the exoskeleton’s adaptation capabilities in a more variable environment. The subject starts from a fixed position and walks to three boxes placed in fixed positions ahead, containing $5$, $10$, or $15$ kg each, as shown in Figure 2 (b). To enhance generalization, we shuffle the positions and contents of the boxes before each test. The subject approaches each box, lifts it individually, and returns to the starting position. This process is repeated three times, ensuring each box is lifted three times. We performed this protocol under three conditions: without the exoskeleton, with static light support (preselected based on subject characteristics), and adaptive pipeline control, which, in its “light support” mode, delivers the same torque as the static light support condition. This setup allows us to compare muscular activity across conditions for each lifting cycle. In this protocol, we allowed participants to lift freely at their preferred speed and style. We designed this experimental condition to both challenge the adaptive control in a more dynamic environment and verify whether the optimal trends observed in standardized tests persist under more natural lifting conditions. As validation metrics, first of all, we keep track during the experiment of the pipeline classification outputs and ground truth. In this context, because latency is an important concern, we only consider correct classifications happening before the lifting cycle starts. We also measure reductions in muscular activity targeting the same muscles of Section III-A, by comparing the no-exoskeleton condition with the static and adaptive control conditions. We collect user feedback using the same discomfort questionnaire described in Section II-A, and we gather user preferences between static and adaptive control. This approach enables us to sample the Validation Representation Functions (VRF) $J_{EMG}(A_{s}^{i},P_{yl}^{i})$, $J_{dsc}(A_{s}^{i},P_{yl}^{i})$, and $J_{prf}(A_{s}^{i},P_{yl}^{i})$ for each static/adaptive assistance (StA/AdA) and payload level pairing, creating a final validation space. Using this space, we assess the performance of adaptive control compared to static control.
@@ IV-A1 EMG-ORF
By analyzing the muscular activation data from the experiments in Section III-A, we derive the EMG-ORF $J_{EMG}(A_{s},P_{yl})$ shown in Figure 4 (I-a). These functions identify optimal performance areas based solely on muscular activity reduction. For instance, reducing back muscle activity consistently requires high assistance, achieving a $6.32\%$ mean and $16.92\%$ peak reduction for light weights, around $9.62\%$ mean and $15.00\%$ peak reduction for medium weights, and a $6.13\%$ mean and a $9.14\%$ peak reduction for heavy weights. However, high assistance for light weights significantly increases peak activation in leg muscles ($15.00\%$), which benefit more from medium or low support. The peak EMG-ORF for leg muscles indicates that transitioning diagonally from light-weight low-assistance to heavy-weight strong-assistance maintains relatively stable peak leg muscle activation while maximizing mean reduction ($5.00\%$ mean reduction for medium assistance with medium weight). Overall, considering all muscles together, medium to strong assistance is preferable for mean reduction, while medium to low for peak. However, given the adverse effect of strong assistance on peak leg muscle activation, a better approach is to use low assistance for light weights and strong assistance for heavy weights. This strategy balances maximizing back muscle activity reduction while minimizing peak activation in the legs.
@@ IV-A2 Discomfort & Preference-ORF
As outlined in Section II-A, we sampled the discomfort function by combining the ratings for each question, adjusting for the negative meaning of discomfort by flipping the ratings of positively worded questions. Using Equation 3, we obtained discomfort samples for different $(A_{s}^{i},P_{yl}^{i})$ pairs. For user preferences, we collected responses on LA/SA preference for LW/HW scenarios. By fitting these samples with GP, we derived continuous formulations of $J_{dsc}$ and $J_{prf}$ within the optimization space, as shown in Figure 4 (I-b). The results reveal that perceived discomfort peaks for medium weights with medium-to-high support. Overall, low assistance is preferable to minimize discomfort, regardless of the payload. Conversely, user preferences indicate a tendency to favor high assistance for heavy weights and low assistance for light weights.
@@ IV-A3 Total-ORF
It’s clear that different ORFs’ contributions can be considered individually to select the best assistance for a given payload for a specific metric optimization. In our case, to maximize acceptability and effectiveness, we combine the normalized versions of EMG-ORF, D-ORF, and P-ORF into a T-ORF $J_{tot}$, considering EMG contributions such as mean and peak reductions in back and overall muscle activity, as shown in Figure 4 (I-c). Results suggest that to optimize EMG reduction, discomfort, and user preference, assistance should be low for light weights and increase with the payload, following a non-linear trend. Based on this, we focus on optimizing “Mean All” activity to minimize muscle activation while avoiding excess leg muscle activation. To determine the optimal assistance for each payload, we compute the partial derivative of $J_{tot}$ with respect to assistance and slice the function along the payload axis. Setting the derivative to zero ($\frac{\partial J_{tot}}{\partial A_{s}}=f(P_{yl},A_{s})=0$) for each chosen $P_{yl}$ yields the optimal $A_{s}^{opt}$ corresponding to each selected $P_{yl}^{i}$, as shown in Figure 5. We subsequently selected an exponential function (see Supplementary Materials for details) to model the relationship between payload and optimal support:
$P_{yl}=a\cdot e^{b\cdot A_{s}^{opt}}+c$ (7)
with $a=0.72$, $b=1.10$, and c=1.20c=-1.20. This function identifies an optimal region in the assistance-payload space, suggesting that if we had a way to continuously estimate payloads, we’d always be able to select the best-performing assistance in terms of our defined metrics. Following these findings, instead of the midpoint between low and high assistance levels we select 73% assistance during validation for medium weights. In the Supplementary Materials, we report the estimated parameters $a$, $b$, and $c$ for the different EMG contributions, together with a detailed justification for the choice of the exponential formulation.